++++ INAF LBC+LUCIFER queue observing, 2011 March 30-31

Observer: R. Speziali, K. Boutsia
Telescope Operator: D. Gonzales
Support Astronomer: O. Kuhn
Instrument: LUCIFER

Summary

Covered sky at twilight
We were closed for 1.5hr due to clouds and later on we closed for high wind
Very difficult to acquire targets because of the clouds. Often the presets were cancelled because of failure to find the guide star
No useful data so far

Data Summary

Project Objects Filter ExpTime Notes
Gallerani SDSSJ1048+4637_ref 210zJHK+Z 1hr we do not see the spectrum - probably due to clouds
Turatto SN2010al Ks 224sec clouds passing by - imaging interrupted - dome closed for clouds

Details

Begin night with Lucifer

12:46 - Taking the darks for the Maiolino HDF_K2 mask that we were not able to take yesterday 0005-0019
  • the previous frames are just test images

02:10 - WFSC collimating * dimm seeing 1.0'' - wfsc seeing estimate 1.0''

02:12 - Slew to Gallerani SDSSJ1048+4637_ref for acq 0020-0023
  • cloudy - waiting for clouds to clear
  • repeat acquisition - we cannot see the target 0024-0025
  • we decide to change target

03:02 - Slew to Turatto SN2010al for acquisition 0028-0031

03:11 -SN2010al - Ks imaging script 0032-0036
  • clouds passing through
  • at the 6th offset we lost guide star and preset was cancelled

03:27 - Repeat preset 0037
  • lost the guide star because of clouds

03:50 - Dome closed for clouds

05:19 - Dome open - Slew to Turatto for acquisition
  • seeing seems to be good so we go to Gallerani

05:26 - Slew to Gallerani SDSSJ1048+4637_ref for acq 0038-0044

06:04 - SG9_spec_science_Z_5.txt 0045-0056
  • there is persistence from the slit in the first image
  • dimm seeing 0.95''
  • although seeing is within the specs, the spectrum is not visible due to the cloudy weather.
  • cloudy - OB must be repeated

07:16 - Take a through-slit to verify that the object is is still centered in the slit 0057
  • we aren't be able to see the object
  • going outside to check the situation by eye - just a few stars are visible.

07:20 - Rotate telescope to measure the wind

07:30 - Dome Closed. Reached wind limits.

-- KonstantinaBoutsia - 31 Mar 2011
Topic revision: r5 - 03 Apr 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback